The sum of the digits of a two digit number is 9. Also, nine times this number is twice the number obtained by reversing the digits of the number. Find the number.
Answer by lwsshak3(11628)

The sum of the digits of a two digit number is 9. Also, nine times this number is twice the number obtained by reversing the digits of the number. Find the number.
**
let u=units digit
let t=tens digit
t+u=9
t=9-u
number=10t+u
..
9(10t+u)=2(10u+t)
90t+9u=20u+2t
11u=88t=88(9-u)=792-88u
99u=792
u=792/99
u=8
t=9-u=1
10t+u=10+8=18
number=18